# Foundation Models For Seismic Data Processing: An Extensive Review

### v1

@@ 3.1.5 Combined Downstream Task Performance
After looking at the individual results by downstream tasks, Figure 6 shows the combined SSIM ($SSIM_{combined}=SSIM_{Demultiple}+SSIM_{Interpolation}+SSIM_{Denoise}$) for each model against the parameters of the FM used as the image encoder. For the combined score the SSIM was chosen, because it has a known upper bound of one, which not only equally weights the different downstream tasks but also gives us a theoretical upper bound of performance. On the left, all different tested models are displayed and on the right, only models which have a combined SSIM above $2.5$ are displayed. This removes all non-hierarchical models, as well as EfficientNet and ViTamin. The best overall model is Swin (V2) pre-trained via self-supervision, followed by Swin (V1) pre-trained via supervision. The third-best model is ConvNeXt (V2) pre-trained via self-supervision. The fourth-best model is CAFormer closely followed by Hiera. Interestingly, CAFormer is performing that well despite being trained via supervision. While the optimal training strategy is task dependent, as mentioned above, across all three downstream tasks self-supervision seems to perform better if available. This could mean, that pre-training CAFormer via self-supervision could improve results significantly, maybe even beyond ConvNeXt (V2).
@@ 3.1.6 Impact of the Pre-Training Dataset
Figure 7 shows that not only the pre-training strategy, but also the dataset and its size, differs. As in Figure 6, all tested models are displayed on the left, and only the ones with a combined SSIM above $2.5$ on the right. In addition, the x-axis of the right plot has a logarithmic scale to better highlight the differences for the smaller datasets, which are otherwise lost with a linear scale. Even though some of the non-hierarchical models were trained on vast datasets, architecture seems to play a more critical role for seismic processing tasks. Nevertheless, even among the hierarchical models there are significant differences in the pre-training dataset sizes that could affect the results. While Swin outperforms ConvNeXt it was also pre-trained on a larger dataset (ImageNet-22k vs. ImageNet-1k) than the other models. The exception are Hiera and the ConvNeXt pre-trained via contrastive learning, both were pre-trained on a significantly larger dataset than the Swin models. More interesting for an SFM, however, is that even models pre-trained on small pre-training datasets yielded good results.
@@ 3.1.7 Impact of Model Size and Inference Time
While only the base model versions are compared in this section, and therefore the number of parameters are similar, two exceptions are EfficientNet and Hiera, which are significantly smaller, see Figure 6. While for EfficientNet this also results in a faster inference time, this is not the case for Hiera, which is only slightly faster than ConvNeXt despite having significantly fewer parameters, as can be seen in Figure 8. Another interesting fact, is that for both ConvNeXt and Swin the V2 versions, which have only minimal architectural changes to accommodate self-supervised pre-training, are slower than their V1 counterparts. The inference times were measured on an A100 using the implementations of Wightman (2019) without torch.compile, a batch size of 512 and gathers of size 224x224. The applicability of these larger FMs is shown by the Swin (V2) inference time of about $2.5$ $\mathrm{ms}$, which translates into around 400 gathers per second. Especially considering, that these inference results are without any optimizations or quantization, which could potentially speed up the model drastically.
3.1.8 Impact of Embedding Space Size
Figure 9: Comparison of the combined performance of the different FMs contrasted against the size of their embedding space. The left plot shows all the tested FMs and the right one only the FMs with a combined SSIM above $2.5$ to further highlight the difference between the well performing models. The well performing models are Swin, ConvNeXt, Hiera, CAFormer and MambaOut. While most of the well performing models have a similar sized embedding space the embedding space of Hiera is significantly smaller.
The final criterion we used to compare these models is the number of pixels in the embedding space, as illustrated in Figure 9. The embedding space $\{E_{s}:E_{s}\in\mathbb{R}^{C_{s}\times H_{s}\times W_{s}}\quad\forall\quad 1\leq s\leq S\}$ is the output of the encoder, as defined in Equation (2). The number of pixels then is the sum over the sizes of each embedding ($\sum_{s=1}^{S}C_{s}\times H_{s}\times W_{s}$). Focusing on the right graph, which again shows only the model with a combined SSIM above $2.5$, we can see that the Swin and ConvNeXt models have the same number of pixels. This is by design, as their channel size was chosen to match a ViT and their down-scaling between stages is identically. However, Hiera has a significantly smaller number of pixels for the same down-scaling factors. This could indicate that smaller channel sizes and therefore smaller models could still lead to good results.
3.1.9 Impact of Model Scaling
In order to ascertain, whether smaller models can still perform well, we compare different model sizes for four archetypes, namely CAFormer, ConvNeXt (V2), Hiera and Swin (V1) on all three seismic processing tasks, see Figure 10. These archetypes were selected because they performed the best across the three tested downstream tasks. For Swin the V1 variant was chosen instead of the V2 variant, because it was available pre-trained in more different model sizes. We opted for the PSNR as the metric instead of the SSIM, as it offers a broader domain, making differentiation between the results of the different models easier.
Figure 10: Performance comparison with scaling model size for the four top performing backbone archetypes on each different downstream task. Plot a) shows the results for the demultiple task, plot b) for the interpolation task and plot c) for the denoising task. As can be seen, smaller Hiera or Swin models still perform well, especially for demultiple and interpolation, where they perform better than significantly larger ConvNeXt models. Additionally, the largest Swin seems to be over-fitted on the pre-training data. The base Hiera performs worse than expected for demultiple and interpolation, but further testing showed that these results are not outliers.
The left plot displays the results for the demultiple task and again confirms our earlier results regarding the ranking for the base model size (around 100M parameters) of Swin $>$ Hiera $>$ CAFormer $>$ ConvNeXt. However, the ranking changes for smaller and larger models, where Hiera outperforms the Swin models. The drop in performance of the Hiera BASE_PLUS version is confounding, but the experiment was repeated an additional five times, even changing the learning rate slightly, and achieved a mean PSNR of $38.96$ with an std of $0.46$, which shows that the original value of $39.36$ was actually a positive outlier. The results for denoising, plotted on the right, also show that the drop in performance is task dependent, because for denoising Hiera scales as excepted. For interpolation the performance of Hiera BASE_PLUS drops again compared to its smaller and larger counterparts, but in contrast to the demultiple results, the Swin models outperform the Hiera models at all scales. The Swin LARGE model appears to be overfitted on the pre-training dataset. However, the smaller Swin models demonstrate notable efficacy, outperforming significantly larger ConvNeXt models. Another outlier is the result of the CAFormer B36 model on denoising, where it exhibits inferior performance compared to its smaller counterparts, despite slightly outperforming them in the other two downstream tasks. To ensure the reliability of these findings, the experiment was repeated five additional times, resulting in an average score of $31.9$ with a standard deviation of $0.14$. While our initial result was slightly worse than the mean performance it remained within expected margins. In conclusion, the findings suggest that smaller Hiera or Swin models could be interesting candidates for a more efficient SFM.
@@ 3.2 Qualitative Results
This section presents qualitative results following the quantitative results in the preceding section. To ensure conciseness, solely the results of the two baseline models (ConvNeXt (A) and Swin (B)) and the overall best non-baseline model (Swin V2 (C)) are displayed. Each of these three models was trained specially for each task, resulting in a total of nine distinct models. Importantly, the following qualitative results are presented to contextualize the quantitative metrics, not to claim improvements over the current state of the art.
@@ 3.2.1 Interpolation
Figure 11: Interpolation results of the two baselines (ConvNeXt (A) and Swin (B)) and the best overall model trained with strategy frozen encoder (Swin V2 (C)) on a gather from the evaluation set. For each model the predicted gather $G^{*}$ as well as the residual $R=G-G^{*}$ is displayed.
Figure 11 depicts the PSNRs of the three models for the interpolation task. The input data for this task is a gather from the evaluation set, which was not observed during training. The first row, consist of this gather $G$ and the decimated gather $D(G)$. The subsequent rows consist of the predicted gather $G^{*}$ and the residual $R=(G-G^{*})\times 10$ of each model. While the predicted gathers ($P^{*}$) appear similar, the differences become more apparent when observing the residual ($R$). While the results of baseline (A) and baseline (B) are similar, the results of model (C) differ significantly. This is confirmed by the calculated PSNRs, which are $39.24$ for (B), $38.05$ for (A) and only $34.96$ for (C). Nonetheless, the results of (C) are impressive considering, that the encoder and therefore the majority of the model (in this case $91\%$ of the total parameters), was trained exclusively on natural images.
@@ 3.2.2 Denoising
Figure 12: Denoising results of the two baselines (ConvNeXt (A) and Swin (B)) and the best overall model trained with strategy frozen encoder (Swin V2 (C)) on a gather from the evaluation set. For each model the predicted gather $G^{*}$ as well as the predicted noise $N^{*}=(G+N)-G^{*}$ is displayed.
Figure 12 depicts the results of the three models for the denoising task. Similar to the interpolation task, a gather from the evaluation set, is used as input. But to better test the generalization ability of these models, uniform noise is added, instead of the Gaussian noise, that was added during training. The first row, consist of a gather $G$, as well as the gather $G+N$. The other rows consist of the predicted gather $G^{*}$ and the predicted noise $N^{*}=(G+N)-G^{*}$ of each model. This time the predicted gathers ($P^{*}$), as well as the predicted noise $N^{*}$ appear similar, across all three models. This is further corroborated by the calculated PSNRs, which are $28.74$ for (B), $27.67$ for (A) and $27.45$ for (C). Impressively, the results of (C) are of comparable quality to the baseline, especially considering again, that the encoder has not seen any seismic data.
@@ 3.2.3 Demultiple
Figure 13: Demultiple results of the two baselines (ConvNeXt (A) and Swin (B)) and the best overall model trained with strategy frozen encoder (Swin V2 (C)) on the North Sea Volve field data. For each model the predicted primaries ($P^{*}$) and the predicted multiples ($M^{*}$) are displayed.
Figure 13 presents the results of these models for the demultiple task on a gather from a post-migration dataset from the North Sea Volve field. The first image is the input gather containing primaries ($P$) and multiples ($M$). The second, third and fourth image depict the predicted primaries ($P^{*}$) by each model. The last three images correspond to the predicted multiples ($M^{*}=(M+P)-P^{*}$), which are derived by subtraction of the predicted primaries from the input gather.
While the predicted primaries ($P^{*}$) appear similar, the differences become more evident when observing the predicted multiples($M^{*}$). The primary dissimilarities of these are in the top third, influenced by the mute and first break. Interestingly, the results from baseline (B) and the overall best model (C) resemble each other more than the results of the two baselines. This indicates that the architecture seems to be more influential than the training strategy when it comes to handling the mute and the first break. The results of (C) are particularly noteworthy, considering that only the decoder was trained on seismic data.
@@ 4.1 Importance of hierarchical models
The most critical finding in this paper is arguably the discrepancy in performance between hierarchical and non-hierarchical models. This finding is corroborated by Goldblum et al. (2023), yet the majority of the FMs in the seismic space utilize a non-hierarchical model, as seen in Sheng et al. (2024), Sansal et al. (2025) and Pham et al. (2025), all of whom employ a ViT. It should be noted, that the results of this investigation are likely more pronounced due to the fact that all the downstream tasks are pixel-level regression tasks. The discrepancy between hierarchical and non-hierarchical models is likely diminished for downstream tasks that do not necessitate pixel level precision, such as classification and segmentation. Nonetheless, we believe that hierarchical models should still exhibit superior performance in these tasks, as indicated by Goldblum et al. (2023), who showed that Swin outperforms ViT in some natural image tasks.
This hypothesis is further substantiated by the observation that the discrepancy between hierarchical and non-hierarchical models is smallest for the denoising task. This can be attributed to the fact that denoising relies less heavily on the higher scale features available through hierarchical models compared to the other two downstream tasks. Noise, being high frequency content, is predominantly present in the higher scale features. Nevertheless, hierarchical models still provide a significant performance gain for denoising, as shown in Figure 5.
Also, the discrepancy between the results of the frozen-encoder strategy and those of the other two is the least significant for denoising, as shown in Figure 4 and Figure 12. This could be attributed to the inherent Gaussian noise in natural images, which the encoder learns through self-supervision without being explicitly trained on denoising.
@@ 4.2 Pre-training strategies and their relevance for an SFM
Another finding less relevant for training an SFM but more interesting from a transfer learning perspective, is that the optimal pre-training strategy differs between downstream tasks. While the reason why self-supervision as pre-training strategy is more optimal for interpolation is clear due to the similarity between both tasks, the reasoning behind the optimal choice for the other two downstream tasks is less clear. These results are similar to those of Goldblum et al. (2023), who showed that the optimal pre-training strategy also differs for natural image tasks. The analysis of pre-training strategies is impeded by the lack of isolated comparison between them. Most pre-training strategies are only available for one model and one dataset. A more fair comparison would be interesting, but would need to be done in the natural image domain. This is because some of the pre-training strategies require a language mapping which is infeasible for seismic data. Additionally, most pre-training strategies require labeled data which is also infeasible for seismic data. Therefore, the only applicable pre-training strategy without synthetic data, for an SFM is self-supervision. Relying on synthetic data semi-supervised learning could also be an alternative, as shown by Cheng et al. (2025).
A notable benefit of self-supervised pre-training on seismic data is its capacity to learn features directly from the data itself. This approach stands in contrast, to the prevailing strategy of training on synthetic data and attempting to enhance it to improve generalization on real data. While Figure 4 demonstrates the comparable or superior performance of fine-tuning a natural image pre-trained model compared to only training on synthetic data, we hypothesize that pre-training on seismic data will yield even better results. A subject that merits further exploration is whether it is preferable to train smaller, more efficient models for significantly different regions or to train a single, large model on all available data.
@@ 4.3 Baseline Performance and Generalization
While the transfer performance of ConvNeXt trained with strategy frozen encoder was significantly worse than Swin, especially when considering different model sizes, the qualitative results for interpolation and denoising showed that the baseline ConvNeXt model that was trained using strategy non-pre-trained encoder was similar in quality if not better than the Swin baseline. We hypothesize that this is due to these examples aligning closely to the training distribution and that Swin models generalize better. This is based on the quantitative results of these models trained with strategy frozen encoder and further corroborated by the qualitative results for demultiple on the North Sea Volve field data.
On the topic of out of distribution generalization, we theorize that the training data used for interpolation and denoising while large enough for in distribution generalization is too small for out of distribution generalization. Therefore, the qualitative results for these two task are shown for in distribution examples. While the interpolation example is completely in distribution, the denoise example is shifted a bit outside distribution by changing the noise distribution but not the underlying seismic distribution. In contrast, the qualitative example for demultiple is from the North Sea Volve field data and therefore out of distribution. Nonetheless, the results are impressive which is in part due to the synthetic training data, which was generated with the explicit intent of wide generalization.
@@ 5 Conclusion
We conducted a comprehensive investigation into various FMs and developed a framework for further testing. Our main focus was on seismic processing, evaluating the performance of the different FMs on three seismic processing tasks: demultiple, interpolation, and denoising. We quantitatively evaluated a diverse range of FMs using synthetic data, aiming to clarify how various model characteristics, such as pre-training techniques and architectural design choices, affect their performance. Additionally, we selected a subset of models for qualitative assessment and presented the demultiple results of real field data. In our comparison of different FMs we also examined the effects of natural image pre-training, with and without additional fine-tuning.
In summary, finding the optimal architecture for seismic processing presents a difficult and multifaceted challenge. However, our research has yielded to some key findings. Firstly, that future SFM, especially if they are intended for seismic processing tasks, should use hierarchical models rather than non-hierarchical ones. Secondly, that Swin seems to be the best overall for models larger than $50M$ parameters, while Hiera may be an alternative option for smaller models. Thirdly, that pre-training, even on natural images, consistently performs as well as or better than no pre-training for all data regimes. Finally, the performance of seismic processing tasks is highly correlated, which supports the creation of an SFM for processing tasks, because a SFMs should collectively maximize its performance on all downstream tasks.
6 Data Availability Statement

### v2

@@ 3.1.3 Impact of the Pre-Training Dataset
Figure 5 reveals variations not only in pre-training strategies but also in the datasets and their sizes. Similar to Figure 4, all tested models are shown on the left, while only those with a combined SSIM above 2.5 are featured on the right. The right plot employs a logarithmic scale on the x-axis to better accentuate differences among smaller datasets, which might be obscured using a linear scale.
Although some non-hierarchical models were trained using extensive datasets, the architecture appears to be more crucial for seismic processing tasks. However, even among hierarchical models, significant differences in pre-training dataset sizes could influence the outcomes. While Swin surpasses ConvNeXt, it also benefits from being pre-trained on a larger dataset (ImageNet-22k vs. ImageNet-1k). Hiera and ConvNeXt trained via contrastive learning are exceptions, having been pre-trained on significantly larger datasets than the Swin models.
More interesting for an SFM is, that models pre-trained on smaller datasets can still achieve impressive results.
@@ 3.1.4 Impact of Model Size and Inference Time
In this section, only the base model versions are compared, resulting in similar parameter counts, with two notable exceptions: EfficientNet and Hiera, which are substantially smaller, as depicted in Figure 4. While EfficientNet benefits from faster inference times, Hiera is only marginally faster than ConvNeXt, despite having significantly fewer parameters, as illustrated in Figure 6. An interesting observation is that for both ConvNeXt and Swin, the V2 versions—which incorporate minimal architectural changes for self-supervised pre-training—are slower than their V1 counterparts.
Inference times were recorded on an A100 using the implementations from Wightman (2019), without torch.compile, with a batch size of 512 and gathers of size 224x224.
The applicability of these larger FMs is demonstrated by the Swin (V2) inference time of approximately 2.5 milliseconds, translating to around 400 gathers per second. Notably, these inference results were obtained without any optimizations or quantization, which both could significantly enhance inference speed.
3.1.5 Impact of natural image pre-training
Figure 7: A comparison of the downstream training strategies for the Swin model family. The observed ranking of these downstream training strategies is fine-tuned encoder $\geq$ non-pre-trained encoder $\geq$ frozen encoder, independently of the downstream task and dataset size.
Figure 7 provides a comparison of the downstream training strategies detailed previously for the Swin model family. The Swin models were selected due to their superior performance, as previously demonstrated. The observed hierarchy of downstream training strategies is fine-tuned encoder $\geq$ non-pre-trained encoder $\geq$ frozen encoder, regardless of downstream task or dataset size. This indicates that pre-training on natural images is consistently comparable to or better than random initialization, especially in low data scenarios, though it holds true even for larger datasets. This suggests that pre-training on natural images may equip the model with capabilities that surpass what can be learned from the specific downstream task alone.
Moreover, the figure corroborates our hypothesis that the performance of the frozen encoder strategy is linked to the other two strategies. Contrary to expectations, the performance gap with the frozen encoder strategy diminishes as more task-specific training data becomes available. Initially, it was thought that pre-training would be beneficial in low-data contexts, as seen with the fine-tuned encoder strategy. The need for more task-specific training data may be due to the domain gap between seismic data and natural images, suggesting that while features learned during pre-training are transferable, the decoder requires sufficient task-specific data to effectively leverage them.
@@ 3.2 Qualitative Results
This section provides qualitative results that complement the quantitative findings presented earlier. For brevity, the focus will be on the outcomes of two baseline models, ConvNeXt (A) and Swin (B), alongside the top-performing non-baseline model, Swin V2 (C). Each model was specifically trained for its respective task, culminating in a total of nine unique configurations. The qualitative results are intended to contextualize the quantitative metrics rather than to assert superiority over the current state of the art.
@@ 3.2.1 Interpolation
Figure 8: Interpolation results of the two baselines (ConvNeXt (A) and Swin (B)) and the best overall model trained with strategy frozen encoder (Swin V2 (C)) on a gather from the evaluation set. For each model the predicted gather $G^{*}$ as well as the residual $R=G-G^{*}$ are displayed.
Figure 8 depicts the results for the interpolation task of each model. The input data comprises a gather from the evaluation set, which was not observed during training. The first row features this gather, denotes as $G$, alongside the decimated gather $D(G)$. Subsequent rows present the predicted gather $G^{*}$ and the residual, calculated as $R=(G-G^{*})\times 10$, of each model. Although the predicted gathers ($P^{*}$) appear similar, the residuals ($R$) reveal more pronounced differences. The results of baseline (A) and baseline (B) show similarities, whereas model (C) displays significant divergence, as confirmed by the PSNRs values of $39.24$ for (B), $38.05$ for (A) and only $34.96$ for (C). Nonetheless, the results of (C) are impressive, considering that the encoder, which in this case accounted for $91\%$ of the total parameters, was trained exclusively on natural images.
@@ 3.2.2 Denoising
Figure 9: Denoising results of the two baselines (ConvNeXt (A) and Swin (B)) and the best overall model trained with strategy frozen encoder (Swin V2 (C)) on a gather from the evaluation set. For each model the predicted gather $G^{*}$ as well as the predicted noise $N^{*}=(G+N)-G^{*}$ are displayed.
Figure 9 presents the results of the three models for the denoising task. To evaluate the models’ generalization capabilities, uniform noise is introduced to the input gather from the evaluation set, differing from the Gaussian noise used during training. The first row displays the gather $G$ and the noisy gather $G+N$. The subsequent rows consist of the predicted gather $G^{*}$ and the estimated noise $N^{*}=(G+N)-G^{*}$ for each model. In this instance, both the predicted gathers ($G^{*}$) and noise estimates ($N^{*}$) appear similar across all three models. This observation is supported by the PSNRs, which are $30.59$ for (B), $31.42$ for (A), and $20.25$ for (C). Model (C)’s results are particularly impressive, as they exhibit quality comparable to the baselines, especially given that the encoder was not exposed to any seismic data.
@@ 3.2.3 Demultiple
Figure 10: Demultiple results of the two baselines (ConvNeXt (A) and Swin (B)) and the best overall model trained with strategy frozen encoder (Swin V2 (C)) on the North Sea Volve field data. For each model, the predicted primaries ($P^{*}$) and the predicted multiples ($M^{*}$) are displayed.
Figure 10 showcases the results for the demultiple task using a gather from a post-migration dataset from the North Sea Volve field. The first image shows the input gather containing primaries ($P$) and multiples ($M$). The following images depict the predicted primaries ($P^{*}$) from each model, while the last three images illustrate the predicted multiples ($M^{*}=(M+P)-P^{*}$), derived by subtracting the predicted primaries from the input gather. While the predicted primaries ($P^{*}$) exhibit similarities, notable differences arise in the predicted multiples ($M^{*}$), particularly in the upper third, influenced by the mute and first break. Interestingly, the results of baseline (B) and model (C) are more alike than those of the two baseline models, suggesting that the architecture may play a more critical role than the training strategy in addressing the mute and first break. The performance of model (C) is especially remarkable, considering that only the decoder was trained on seismic data.
@@ 4.1 Importance of hierarchical models
The most critical finding in this paper is arguably the discrepancy in performance between hierarchical and non-hierarchical models. However, the majority of the FMs in the seismic space utilize a non-hierarchical model, as seen in Sheng et al. (2024), Sansal et al. (2025) and Pham et al. (2025), all of whom employ a ViT. It should be noted, that the results of this investigation are likely more pronounced due to the fact that all the downstream tasks are pixel-level regression tasks. The discrepancy between hierarchical and non-hierarchical models is likely diminished for downstream tasks that do not necessitate similar precision, such as salt body classification or fault detection. Nonetheless, we believe that hierarchical models should still exhibit superior performance in these tasks, as corroborated by Goldblum et al. (2023).
@@ 4.2 Pre-training strategies and their relevance for an SFM
The analysis of pre-training strategies is impeded by the lack of isolated comparison between them. Most pre-training strategies are only available for one model and one dataset. A more fair comparison would be interesting, but would need to be done in the natural image domain. This is because some of the pre-training strategies require a language mapping which is infeasible for seismic data. Additionally, most pre-training strategies require labeled data which is also infeasible for seismic data. Therefore, the only applicable pre-training strategy without synthetic data, for an SFM is self-supervision. Relying on synthetic data semi-supervised learning could also be an alternative, as shown by Cheng et al. (2025).
A notable benefit of self-supervised pre-training on seismic data is its capacity to learn features directly from the data itself. This approach stands in contrast to the prevailing strategy of training on synthetic data and attempting to enhance it to improve generalization on real data. While Figure 7 demonstrates the comparable or superior performance of fine-tuning a natural image pre-trained model compared to only training on synthetic data, we hypothesize that pre-training on seismic data will yield even better results. A subject that merits further exploration is whether it is preferable to train smaller, more efficient models for significantly different regions or to train a single, large model on all available data.
@@ 4.3 Baseline Performance and Generalization
While the transfer performance of ConvNeXt trained with strategy frozen encoder was significantly worse than Swin, the baseline ConvNeXt model that was trained using strategy non-pre-trained encoder performed better than the Swin baseline. We hypothesize that this is due to these examples aligning closely to the training distribution and that Swin models generalize better. This is based on the quantitative results of these models trained with strategy frozen encoder and further corroborated by the qualitative results for demultiple on the North Sea Volve field data.
On the topic of out-of-distribution generalization, we theorize that the training data used for interpolation and denoising while large enough for in-distribution generalization is too small for out-of-distribution generalization. Therefore, the qualitative results for these two task are shown for in-distribution examples. While the interpolation example is completely in-distribution, the denoise example is shifted a bit outside distribution by changing the noise distribution but not the underlying seismic distribution. In contrast, the qualitative example for demultiple is from the North Sea Volve field data and therefore out-of-distribution.
@@ 5 Conclusion
We carried out an extensive study on various FMs and developed a framework for further evaluation. Our primary focus was on seismic processing, assessing the performance of different FMs across three seismic processing tasks: demultiple, interpolation, and denoising. We quantitatively assessed a wide array of FMs using synthetic data to elucidate how different model attributes, such as pre-training strategies and architectural design choices, influence their performance. Furthermore, we selected a subset of models for qualitative analysis and presented demultiple results using real field data. Our comparison of different FMs also explored the impact of pre-training on natural images, with and without additional fine-tuning.
Identifying the optimal architecture for seismic processing is a complex and multifaceted challenge. However, our research has led to several key insights. First, future SFMs, particularly those intended for seismic processing tasks, should favor hierarchical models over non-hierarchical ones. Second, Swin appears to be the best overall model, followed by ConvNeXt. Lastly, pre-training on natural images consistently performs as well as or better than training from scratch across all data scenarios.
6 Acknowledgments
The authors would like to acknowledge the members of the Fraunhofer ITWM DLSeis Consortium (http://dlseis.org) for their financial support. We are grateful to Equinor and Volve Licence partners for releasing The North Sea Volve Dataset under an Equinor Open Data Licence. We are also grateful to Hemang Shah and BP Exploration Operation Company Limited ("BP") for providing the 2007 BP Anisotropic Velocity Benchmark dataset.
7 Data and Materials availability